# Universal activated aging and weak ergodicity breaking in spin and structural glasses

### v4

@@ 3. Estimate of the onset temperature $T_{\rm onset}$
Figure S4: Comparison between the MC aging function (points) and Eq. (S24) (lines), at $T=0.75$. No fitting parameters are used for the lines.
@@
S2 Thermodynamics of the finite-size random energy model: a tree-expansion theory
S2.1 Probability distributions
@@
Our aim is to obtain statistical properties of the energy landscape for the finite-sized REM with NN spins, where NN is sufficiently small so that the system is far from the thermodynamic limit ($N\to\infty$). We assume that each basin only has one nondegenerate local minimum, and that basins are independent. The probability distribution of the local minimum $E_{\rm lm}$ is,
@@
where ${L}(E)\equiv\int_{E}^{\infty}dE\rho(E)$ is a complementary cumulative distribution function, and $\rho(E)$ is the probability distribution function of the configuration energy $E$. The above expression requires that the NN direct neighbors of the local minimum, which are related to the local minimum by a single spin flip, have an energy $E$ higher than $E_{\rm lm}$. The factor $N+1$ comes from the $N+1$ choices of the local minimum among the $N+1$ configurations. For the G-REM, $\rho(E)=\rho_{\rm Gauss}(E)$ given by Eq. (7), which results in ${L}(E)={L}_{\rm Gauss}(E)$, where
@@
S2.1.1 The first order: $\Omega=1$
When $\Omega=1$, as shown in Fig. S5, the basin contains only the local minimum without any other configurations. By definition, the saddle point should have the lowest energy among the NN neighbors of the local minimum – the probability of this condition is ${L}^{N-1}(E_{{\rm sp}})$. The saddle point also has NN neighbors, including the local minimum and N1N-1 other neighbors. The energies of the other N1N-1 neighbors cannot be all higher than $E_{{\rm sp}}$ – otherwise the saddle point would be a configuration inside the basin (not a saddle point). This condition imposes a constraint given by $1-{L}^{N-1}(E_{{\rm sp}})$. Putting these considerations together, we can write the first-order joint distribution as
@@
where the extra factor $(N+1)N$ comes from the permutation of choosing a local minimum from the $N+1$ configurations and a saddle point from the rest of the NN configurations.
Figure S5: Graphic representation of the first-order joint distribution Eq. (S31). Each node represents a configuration, and the height of the node represents the configuration energy $E$. Any pair of configurations connected by a link are related by a single-spin flip. The saddle point energy $E_{\rm sp}$ is marked by the dotted line. For $\Omega=1$, there is only one configuration in the basin ($E<E_{\rm sp}$), which is the local minimum. The number of nodes in the circled cluster is indicated (in this case, both clusters have N1N-1 nodes). For simplification, only $N=4$ nodes are shown.
S2.1.2 The second order: $\Omega=2$
Figure S6: Graphic representation of the second-order joint distribution Eq. (S32).
@@
S2.1.3 The third order: $\Omega=3$
Figure S7: Graphic representation of the third-order joint distribution Eq. (S33).
@@
Figure S8: Comparison between the tree-expansion theory (lines) and the barrier-tree method (points), for the G-REM and E-REM with $N=24$ spins. (a) Probability distribution function (pdf) of the local minimum energy. The line represents Eq. (S25). (b) Pdf of the saddle point energy. (c,d) Pdf of the barrier energy distribution. In (b-d), the red line represents the tree-expansion theoretical results up to the first-order ($\Omega=1$), and the black line represents the results up to the third-order results ($\Omega=3$).
S2.1.4 Verification with the numerical barrier-tree algorithm
@@
Figure S9: Probability $g_{N}(\Omega)$ of having $\Omega$ configurations in a basin. Lines are tree-expansion theory results Eq. (S35), and points are numerical data obtained by the barrier-tree algorithm.
S2.2 Internal entropy
@@
Figure S10: Probability distribution function of the internal entropy for the G-REM. The data points are $g^{\rm BT}(S)$ obtained by the numerical barrier-tree method, and the solid line is Eq. (S36). (inset) Average numerical entropy $\langle S\rangle$ as a function of NN, where the dashed line is $S_{\infty}\approx 0.212439$ (Eq. S37).
@@
This result is consistent with the numerical data obtained by the barrier-tree method (see Fig.  S10). The inset of Fig. S10 shows that the mean internal entropy $\langle S\rangle$ converges to a value $S_{\infty}\approx 0.21$ in the large-NN limit. Then the average number of configurations in a basin is $\Omega\approx e^{0.21}\approx 1.2$, i.e., each basin contains approximately one configuration in average. Thus in the spin glass phase, the mean internal entropy per spin vanishes ($\langle S\rangle/N\to 0$) in the thermodynamic limit, consistent with thermodynamic theory of the REM [33].
S2.3 Number of basins
We assume that each basin contains one local minimum. Two local minima cannot be directly related via a single-spin flip – otherwise the two basins cannot be separated. Thus, in average, we can assume that each basin occupies $N+1$ configurations, i.e., one local minimum and its NN neighbours. The total number of configurations for a NN-spin REM is $2^{N}$. Therefore, the total number of basins is,
@@
Figure S11: Number of basins $N_{\rm b}$ as a function of the system size NN. The line represents Eq. (S38), and the points are data obtained by the numerical barrier-tree algorithm.

### v5

@@ Acknowledgments
FIG. S5: Comparison between the MC aging function (points) and Eq. (S24) (lines), at $T=0.75$. No fitting parameters are used for the lines.
S2 Additional simulation data for the random energy model
S2.1 Energy trajectory
We show an example of the time trajectory of the configuration energy $E(t_{\rm w})$ for the G-REM (Fig. S6a), and the corresponding trajectory of the basin (saddle point) energy $\tilde{E}(t_{\rm w})$ (Fig. S7b).
FIG. S6: An energy trajectory obtained by the MC simulation of the G-REM. Data are obtained at $T=0.75$ with $N=20$. (a) An example of configuration trajectory, where $E(t_{\rm w})$ is the energy of the configuration at $t_{\rm w}$. (b) Corresponding basin trajectory, where $\tilde{E}(t_{\rm w})$ is the energy of the basin or the saddle point at $t_{\rm w}$. With this, a sequence of hopping events with the barrier energy $\Delta E_{i_{k}}$ and the hopping time $\tau_{k}$ are identified.
S2.2 Test of the renewal mechanism
Occasionally, after leaving a basin, the system jumps back to the same basin. Following the previous study [38], the two events are merged in such cases. The effects of return jumps are revealed in Fig. S7(a,b) by the conditional probability $p(\tau_{2}|\tau_{1})$, which is the distribution of the next-event hopping time $\tau_{2}$ when the current-event hopping time is $\tau_{1}$. Only when the return jumps are merged, the $p(\tau_{2}|\tau_{1})$ is independent of $\tau_{1}$, which means that the two consecutive jumps are independent. The independence of consecutive jumps (the renewal mechanism) is a basic assumption in the trap model. In order to be consistent with the renewal mechanism, we employ the no-return treatment.
FIG. S7: Test of the renewal mechanism in the G-REM. Data are obtained by MC simulations at $T=0.75$ with $N=20$. (a) The hopping time probability distribution function (pdf) $\psi(\tau)$ and the conditional distribution $\psi(\tau_{2}|\tau_{1})$ measured in MC simulations, (b) without and (d) with return hops; $\psi^{\rm BT}(\tau)$ is converted from the static distribution $p^{\rm BT}_{\rm G-REM}(\Delta E)$ using the Arrhenius law.
S2.3 Verification of the Arrhenius law
$\tau(\Delta E)=A_{0}\exp(\Delta E/T).$ (S25)
We find that the Arrhenius law is consistent with the $\bar{\tau}(\Delta E)$ data ($\bar{\tau}$ is the mean hopping time for the given $\Delta E$) obtained by MC simulations (Fig. S8a), with $A_{0}\sim O(1)$ as shown in Fig. S8b-inset (the time unit is $N$).
When $N$ is finite, for a given $\Delta E$, $\tau$ follows a distribution
$\Psi(\tau|\Delta E)=(1-\eta)^{\tau-1}\eta\approx e^{-\tau\eta}\eta,$ (S26)
where $\eta=\exp(-\Delta E/T)$. To derive $\Psi(\tau|\Delta E)$, consider a discretized time $\tau$: if a system is trapped in a basin with a $\Delta E$-barrier for $\tau$ steps, it should remain in the basin for $\tau-1$ steps with a probability $1-\eta$ at each step, and jumps out the basin at the final step with a probability $\eta$. Equation (S26) is also verified by the simulation data (see Fig. S8b). Note that the mean hopping time of Eq. (S26), $\bar{\tau}(\Delta E)=\int_{0}^{\infty}\tau\Psi(\tau|\Delta E)d\tau$, consistently recovers the Arrhenius law Eq. (S25). In short, the Arrhenius law is verified by our MC data.
FIG. S8: Arrhenius law. (a) Correspondence between $\tau$ and $\Delta E$ in MC simulations of the G-REM at $T=0.75$ with $N=20$, where the color represents the logarithm of the joint probability $\ln\Psi(\tau,\Delta E)$ (see the color bar). (b) The conditional probability $\Psi(\tau|\Delta E=10)$ multiplied by $\tau$; the solid line follows Eq. (S26). (b-inset) $A_{0}$ vs NN.
@@
S3 Thermodynamics of the finite-size random energy model: a tree-expansion theory
S3.1 Probability distributions
@@
Our aim is to obtain statistical properties of the energy landscape for the finite-sized REM with $N$ spins, where $N$ is sufficiently small so that the system is far from the thermodynamic limit ($N\to\infty$). We assume that each basin only has one nondegenerate local minimum, and that basins are independent. The probability distribution of the local minimum $E_{\rm lm}$ is,
@@
where ${L}(E)\equiv\int_{E}^{\infty}dE\rho(E)$ is a complementary cumulative distribution function, and $\rho(E)$ is the probability distribution function of the configuration energy $E$. The above expression requires that the $N$ direct neighbors of the local minimum, which are related to the local minimum by a single spin flip, have an energy $E$ higher than $E_{\rm lm}$. The factor $N+1$ comes from the $N+1$ choices of the local minimum among the $N+1$ configurations. For the G-REM, $\rho(E)=\rho_{\rm Gauss}(E)$ given by Eq. (7), which results in ${L}(E)={L}_{\rm Gauss}(E)$, where
@@
S3.1.1 The first order: $\Omega=1$
When $\Omega=1$, as shown in Fig. S9, the basin contains only the local minimum without any other configurations. By definition, the saddle point should have the lowest energy among the $N$ neighbors of the local minimum – the probability of this condition is ${L}^{N-1}(E_{{\rm sp}})$. The saddle point also has $N$ neighbors, including the local minimum and N1N-1 other neighbors. The energies of the other N1N-1 neighbors cannot be all higher than $E_{{\rm sp}}$ – otherwise the saddle point would be a configuration inside the basin (not a saddle point). This condition imposes a constraint given by $1-{L}^{N-1}(E_{{\rm sp}})$. Putting these considerations together, we can write the first-order joint distribution as
@@
where the extra factor $(N+1)N$ comes from the permutation of choosing a local minimum from the $N+1$ configurations and a saddle point from the rest of the $N$ configurations.
FIG. S9: Graphic representation of the first-order joint distribution Eq. (S33). Each node represents a configuration, and the height of the node represents the configuration energy $E$. Any pair of configurations connected by a link are related by a single-spin flip. The saddle point energy $E_{\rm sp}$ is marked by the dotted line. For $\Omega=1$, there is only one configuration in the basin ($E<E_{\rm sp}$), which is the local minimum. The number of nodes in the circled cluster is indicated (in this case, both clusters have N1N-1 nodes). For simplification, only $N=4$ nodes are shown.
S3.1.2 The second order: $\Omega=2$
FIG. S10: Graphic representation of the second-order joint distribution Eq. (S34).
@@
S3.1.3 The third order: $\Omega=3$
FIG. S11: Graphic representation of the third-order joint distribution Eq. (S35).
@@
FIG. S12: Comparison between the tree-expansion theory (lines) and the barrier-tree method (points), for the G-REM and E-REM with $N=24$ spins. (a) Probability distribution function (pdf) of the local minimum energy. The line represents Eq. (S27). (b) Pdf of the saddle point energy. (c,d) Pdf of the barrier energy distribution. In (b-d), the red line represents the tree-expansion theoretical results up to the first-order ($\Omega=1$), and the black line represents the results up to the third-order results ($\Omega=3$).
S3.1.4 Verification with the numerical barrier-tree algorithm
@@
FIG. S13: Probability $g_{N}(\Omega)$ of having $\Omega$ configurations in a basin. Lines are tree-expansion theory results Eq. (S37), and points are numerical data obtained by the barrier-tree algorithm.
S3.2 Internal entropy
@@
FIG. S14: Probability distribution function of the internal entropy for the G-REM. The data points are $g^{\rm BT}(S)$ obtained by the numerical barrier-tree method, and the solid line is Eq. (S38). (inset) Average numerical entropy $\langle S\rangle$ as a function of $N$, where the dashed line is $S_{\infty}\approx 0.212439$ (Eq. S39).
@@
This result is consistent with the numerical data obtained by the barrier-tree method (see Fig.  S14). The inset of Fig. S14 shows that the mean internal entropy $\langle S\rangle$ converges to a value $S_{\infty}\approx 0.21$ in the large-$N$ limit. Then the average number of configurations in a basin is $\Omega\approx e^{0.21}\approx 1.2$, i.e., each basin contains approximately one configuration in average. Thus in the spin glass phase, the mean internal entropy per spin vanishes ($\langle S\rangle/N\to 0$) in the thermodynamic limit, consistent with thermodynamic theory of the REM [33].
S3.3 Number of basins
We assume that each basin contains one local minimum. Two local minima cannot be directly related via a single-spin flip – otherwise the two basins cannot be separated. Thus, in average, we can assume that each basin occupies $N+1$ configurations, i.e., one local minimum and its $N$ neighbours. The total number of configurations for a $N$-spin REM is $2^{N}$. Therefore, the total number of basins is,
@@
FIG. S15: Number of basins $N_{\rm b}$ as a function of the system size $N$. The line represents Eq. (S40), and the points are data obtained by the numerical barrier-tree algorithm.
S4 Additional simulation data for structural glasses
S4.1 Comparison of different simulation algorithms for the WCA model.
We have checked that the aging functions $F_{w}(t_{\rm w})$ obtained by different simulation methods are consistent with each other, apart from the short-time dynamics that are protocol-dependent (see Fig. S16a). Four different simulation methods are compared. (i) MD simulations in the isokinetic ensemble. (ii) MD simulations in the microcanonical (NVE) ensemble with the kinetic energy rescaled to the target value every 50 MD steps (0.1 LJ time units). (iii) MD simulations in the canonical (NVT) ensemble using the Nose-Hoover chain thermostat with a damping parameter of 0.1 LJ time units [85]. (iv) Standard MC simulations in the canonical ensemble. In the MC simulation, we attempt to translate a particle by a displacement vector randomly drawn in a sphere of a radius $5\times 10^{-2}\sigma_{\rm AA}$. The move is accepted according to the Metropolis acceptance rule. In Fig. S16, the MC steps are multiplied by $6.0\times 10^{-4}$ to match the long-time dynamics with those obtained by MD simulations. Further comparison is carried out for the non-equilibrium two-time mean-squared displacement (MSD), $\delta r^{2}(t_{\rm w},t_{\rm w}+t)\equiv\frac{1}{N_{\rm A}}\langle\sum_{i=1}^{N_{\rm A}}|{\bf r}_{i}(t_{\rm w}+t)-{\bf r}_{i}(t_{\rm w})|^{2}\rangle$ (see Fig. S16b). These tests validate the isokinetic ensemble adapted in our MD simulations, which is used to generate the results reported in the current study.
FIG. S16: Comparison of the results obtained by different MD and MC algorithms for the WCA model. (a) Aging function $F_{w}(t_{\rm w})$ and (b) non-equilibrium two-time MSD $\delta r^{2}(t_{\rm w},t_{\rm w}+t)$ for $N=1000$ particles at $T=0.25$. In panel (b), $t_{\rm w}\approx 1.0~,16.0,~135,~1.15\times 10^{3},~9.65\times 10^{3},~4.21\times 10^{4}$ from left to right.
S4.2 Absence of crystallization and phase separation during aging in the WCA model.
The growth of crystals is marked by the emergence of Bragg peaks in the measured structure factor, $S(q)$, which are absent in the liquid state [89]. Therefore, a constant $S(q)$ is a standard criterion for ruling out crystallization in amorphous systems [50]. During aging of the WCA system, we observe negligible change in $S(q)$ for both total and partial static structure factor (Fig. S17a-b), which suggests that crystallization is not occurring. This finding is consistent with the fact that the nucleation time, $\tau_{X}$, typically exceeds the $\alpha$-relaxation time, $\tau_{\rm eq}$, by orders of magnitude [90]; since aging takes place prior to equilibration, crystallization is not expected on this timescale.
A more sensitive probe of long-range order is $S_{0}=S(q\to 0)$, obtained by extrapolating the short-wavelength behavior of $S(q)$ (Fig. S17a-b, inset). Within the Ornstein-Zernike theory, $S_{0}$ is proportional to the square of the correlation length, $\xi_{\rm OZ}^{2}$, which characterizes positional pair correlations [91]. Our data show that $S_{0}$ initially decreases weakly with waiting time, $t_{\rm w}$, during aging (Fig. S17c-d). Only after $t_{\rm w}\approx 10^{5}$, does $S_{0}$ begin to increase. This late-stage growth implies the development of long-range order, possibly due to subtle fractionation or phase separation. However, these effects are minimal, as evidenced by the similar behavior of the full $S(q)$ and the partial $S^{AA}(q)$ for large (type A) particles. Moreover, the finite-size effects are insignificant.
Crucially, these potential fractionation effects manifest at $t_{\rm w}$ values far beyond the aging regime central to this study. This is clear from two observations. First, the aging relation $\tau_{\rm neq}\sim t_{\rm w}$ holds only for $t_{\rm w}<10^{3}$ (Fig. 1b). Second, the parameter $k$, from which the aging length $\xi_{\rm ag}$ is derived, is extracted from the data with $t_{\rm w}<10^{3}$ (see Fig. 5b). In Fig. S17c-d, we use arrows to indicate the time at which the logarithmic decay, $F_{w}(t_{\rm w})\sim-k\ln t_{\rm w}$, ends. It can be seen that the growth of $S_{0}$ emerges after this logarithmic regime. We can therefore safely conclude that crystallization and phase separation are negligible within the aging dynamics under investigation.
FIG. S17: Static structural factor $S(q)$ for the WCA model. Data are obtained at $T=0.25$, quenched from $T_{\rm eq}=5$. The structure factor for (a) all particles and (b) large particles, at a few different $t_{\rm w}$. Inset is the close-up for small $q$. Estimates for $S_{0}$ and $S_{0}^{AA}$ are obtained by taking the mean of the five data points corresponding to the smallest $q$. (c) $S_{0}$ and (d) $S_{0}^{AA}$ as functions of $t_{\rm w}$, for a few different NN. The arrows indicate the end of the logarithmic decay of $F_{w}(t_{\rm w})$.
S4.3 Estimate of the onset temperature for amorphous silica
The onset temperature $T_{\rm onset}$ is defined as the characteristic temperature below which glassy dynamics appear. For amorphous silica, $T_{\rm onset}$ has not yet been reported in the literature. In this study two independent methods are employed to estimate $T_{\rm onset}$. The first method is based on the $T$-dependent behavior of the $\alpha$-relaxation time $\tau_{\rm eq}(T)$ in supercooled liquids. One defines $T_{\rm onset}$ at the point where $\tau_{\rm eq}(T)$ departs from the high-$T$ Arrhenius behavior, $\tau_{\rm eq}(T)\sim\exp(E_{\infty}/T)$. Applying this criterion to our simulation data gives $T_{\rm onset}\approx 5800$ K (see Fig. S18a). Note that for amorphous silica, $\tau_{\rm eq}(T)$ is also Arrhenius at low $T$ (with a constant, but larger activation energy than $E_{\infty}$), consistent with the well-know fact that amorphous silica forms strong glasses.
The second method is based on the $T$-dependent behavior of the potential energy $E_{\rm IS}(T)$ of inherent structures [92]. The inherent structures at zero temperature are obtained by minimizing the energy of configurations initially equilibrated at $T$ using the FIRE algorithm [93]. The high-$T$ and low-$T$ data of $E_{\rm IS}(T)$ are fitted by two linear functions, whose intersection defines $T_{\rm onset}\approx 5800$ K (see Fig. S18b). The above two methods give a consistent $T_{\rm onset}$ within the numerical uncertainty.
FIG. S18: Onset temperature $T_{\rm onset}$ in the amorphous silica model. (a) $\alpha$-relaxation time $\tau_{\rm eq}$ as a function of $1/T$. The high-$T$ data are fitted to the Arrhenius law (line), and the departing point indicates $T_{\rm onset}$. (b) Inherent structure energy $E_{\rm IS}$ as a function of $T$. The two lines represent linear fitting to low-$T$ and high-$T$ data, whose intersection defines $T_{\rm onset}$. Data are obtained for $N=1536$.
S4.4 Nonergodicity order parameter
The nonergodicity parameter $f$ is defined by $f=\lim_{t\to\infty}F_{\rm eq}(t)$, where $F_{\rm eq}(t)$ is the equilibrium incoherent scattering function. For the aging data, $f$ corresponds to the plateau value of $F(t_{\rm w},t_{\rm w}+t)$ by sending $t_{\rm w}\to\infty$ first (equilibrium limit) and then $t\to\infty$ (large-time limit); see Fig. 1c. Figure S19 shows the data of $f$ obtained for both WCA and amorphous silica models. Interestingly, the data nearly collapse if plotted as functions of the rescaled temperature $T/T_{\rm WEB}$.
FIG. S19: Nonergodicity parameter $f$ with respect to the rescaled temperature $T/T_{\rm WEB}$, for WCA and amorphous silica models.
S4.5 Exponent $\beta$
According to our theory, the aging function follows logarithmic behavior in the intermediate-time regime, $F_{w}(t_{\rm w})\sim-k\ln t_{\rm w}$. For mean-field models (e.g., REMs), $k\sim 1/N$ (see Eq. 5). For finite-dimensional models, such as the WCA and amorphous silica models, we find that $k\sim N^{-\beta}$ as shown in Fig. 5(c, f) (in the WCA model, $k$ becomes a constant when $N>n$). The $T$-dependent exponent $\beta$ is reported in Fig. S20. For both models, $\beta\approx 0.1-0.4$, which is considerably smaller than the mean-field value $\beta=1$. It is unclear if this discrepancy comes from finite-dimensional effects or pre-asymptotic effects due to insufficiently large systems.
FIG. S20: Exponent $\beta$ as a function of $\hat{x}=T/T_{\rm WEB}$ in WCA and amorphous silica models.